SOME REMARKS ON REAL AND COMPLEX OUTPUT FEEDBACK

JOACHIM ROSENTHAL FRANK SOTTILE

Abstract. We provide some new necessary and sufficient conditions which guarantee arbitrary pole placement of a particular linear system over the complex numbers. We exhibit a non-trivial real linear system which is not controllable by real static output feedback and discuss a conjecture from algebraic geometry concerning the existence of real linear systems for which all static feedback laws are real.